The equilibrium constant for the formation of Cu(CN)42- is 2.0 x 1030. Calculate the value of pCu2+, i.e. -log[Cu2+], if 2.48 g of CuCl2 is dissolved in 1.000 L of a 0.937 M NaCN solution. The addition of CuCl2 does not affect the volume (the final volume is always 1.000 L).


1
Expert's answer
2021-11-24T12:24:17-0500

CuCl2 +4NaCN = Na2[Cu(CN)4] + 2NaCl

Cu(CN)42- = Cu2+ + 4CN-

K = [Cu2+][ CN-]4 / [Cu(CN)42-] = 2×1030

Molecular weight of CuCl2 is 134.45 g/mol

Number of mol CuCl2 is = 2.76/134.45 mol =0.0205 mol

Number of mol NaCN is = 0.848 × 1000 = 848 mol 

 [Cu(CN)42-] = 0.0205 mol and CM = 0.0205/1000 = 2.05 × 10-5

Cu(CN)42- = Cu2+ + 4CN-     [Cu2+ = S, CN- =S]

K = S ×4S4 = 4S5

K×[Cu(CN)42-] = 4S5

S =[(2×1030) × (2.05 × 10-5)/4]1/5 = 100495.0737

p[Cu2+] or -log[Cu2+] = -log(100495.0737) = -5.002 (Answer)